OpenStudy (anonymous):

Find the area of the region that lies inside both curves: r=3sin(2θ) r=3sin(θ)

OpenStudy (anonymous):

Find the point of intersection and use that as your boundary points when doing the integral.

OpenStudy (anonymous):

I did! Can you work it out and see what you get?

OpenStudy (anonymous):

\[\int\limits_{a}^{b} (1/2)r^2 dr\] is the formula

OpenStudy (anonymous):

like i totally forgot how to graph in polr coordinates

OpenStudy (anonymous):

but i cant get it to work even trying different bounds

OpenStudy (apoorvk):

lol, i didnt read this was polar coordinates. i don't have much idea, but lets see.. i had drawn a complete graph, and was just about to post it stupidly, when i happened to see samjordon's comment above.

OpenStudy (anonymous):

Helppp meeee somoneeeee

OpenStudy (anonymous):

i think i am giving up cuz its 1 30 am and i didnt finish my own work and i will be up till 4 in the morning. sorry

OpenStudy (anonymous):

@KingGeorge and @eliassaab can u guys come help out

OpenStudy (kinggeorge):

Well, let's find the intersection points.\[3\sin(\theta)=3\sin(2\theta)\]So \[\sin(\theta)=\sin(2\theta)\]The obvious solution is \(\theta=0\). The other solutions are at \(\theta=\pi/3, \;\;\pi, \;\;5\pi/3\)

OpenStudy (anonymous):

We are solving for the area

OpenStudy (anonymous):

Remember it is in polar

OpenStudy (kinggeorge):

Those give us the bounds of integration. Give me a minute to make sure we're integrating the correct things here.

OpenStudy (anonymous):

ok

OpenStudy (kinggeorge):

After I've graphed it it looks a bit wonky. it looks sort of like And we want to find the area of the shaded region.